OpenStudy (anonymous):

Can someone help and explain two questions to me? o: ((Square roots, 8th grade))

OpenStudy (anonymous):

sure post the question.

OpenStudy (anonymous):

i dont really understand this one ;-;

OpenStudy (surana):

Wait, do you want help with square roots or explaining what they are?

OpenStudy (anonymous):

i need help with the exponent square roots, like the question i posted

OpenStudy (surana):

I hate to admit it, but this isn't something I can do readily. Sorry. You could try using the Symbolab Algebra Calculator, I've used it before and it's useful.

OpenStudy (anonymous):

okay, but thank you anyways :3

OpenStudy (isaiah.feynman):

\[a^{\frac{ m }{ n }} = \sqrt[n]{a^{m}}\]

OpenStudy (anonymous):

ohh, okay, i didn't think my first answer was right, but i had gotten A. x^7, which i think is maybe right?

OpenStudy (isaiah.feynman):

For your question, take n = 5 and m = 14.

OpenStudy (anonymous):

hm, okay, so i put is like: \[a = 14/5\]

OpenStudy (anonymous):

okay, i got the answer, but thank you for helping me with the formula.

OpenStudy (isaiah.feynman):

Good stuff.